A research conducted on a certain company last year revealed that 25% of the employees prefer drinking milk tea than coffee during break time The company has decided to give free milk tea during break time in a recent study out of 100 randomly sampled employees 28% said that they would rather drink milk tea than coffee. What is the z-value?

1
Expert's answer
2022-06-16T06:24:23-0400

The z-statistic is computed as follows:



"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.28-0.25}{\sqrt{\dfrac{0.25(1-0.25)}{100}}}=0.6928"